# Hyp-OW: Exploiting Hierarchical Structure Learning with Hyperbolic Distance Enhances Open World Object Detection

### v1

@@ Abstract
Open World Object Detection (OWOD) is a challenging and realistic task that extends beyond the scope of standard Object Detection task. It involves detecting both known and unknown objects while integrating learned knowledge for future tasks. However, the level of ’unknownness’ varies significantly depending on the context. For example, a tree is typically considered part of the background in a self-driving scene, but it may be significant in a household context. We argue that this external or contextual information should already be embedded within the known classes. In other words, there should be a semantic or latent structure relationship between the known and unknown items to be discovered. Motivated by this observation, we propose Hyp-OW, a method that learns and models hierarchical representation of known items through a SuperClass Regularizer. Leveraging this learned representation allows us to effectively detect unknown objects using a Similarity Distance-based Relabeling module. Extensive experiments on benchmark datasets demonstrate the effectiveness of Hyp-OW achieving improvement in both known and unknown detection (up to 6 points). These findings are particularly pronounced in our newly designed benchmark, where a strong hierarchical structure exists between known and unknown objects.
1 Introduction
@@
Advances in Object Detection (OD) have unlocked a plethora of practical applications such as robotics  Zhou et al. (2022), self-driving cars  Balasubramaniam and Pasricha (2022), manufacturing  Malburg et al. (2021), and medical analysis  Yang and Yu (2021). Recent breakthroughs in attention-based neural network architecture, such as Deformable Transformers  Zhu et al. (2021), have yielded impressive performance in these settings. However, most of these approaches assume a fixed number of classes (closed-world assumption), which is rare in reality. Continual Object Detection  Menezes et al. (2023) takes a step further by incrementally adding new classes, resulting in a distribution shift in the input and the well-known phenomenon of catastrophic forgetting(Kirkpatrick et al., 2017; Doan et al., 2021) where the network forgets previously learned knowledge. Open World (OW) Bendale and Boult (2015) takes these assumptions even further, introducing the detection and integration of newly discovered classes.
@@
While the seminal work by  Bendale and Boult (2015) introduced the Open World (OW) framework, further advancements by  Joseph et al. (2021) extended it in two key aspects: the detection task and continual learning. However, a significant challenge within this framework lies in the absence of annotations for unknown objects, leading to biases towards known labels and potential confusion between unknown items and the background. This bias significantly impedes the accurate identification of unknown objects and presents a major hurdle in the detection process.
We can summarize previous attempts to solve this problem into three main categories. The first category includes works such as Joseph et al. (2021); Gupta et al. (2022); Zohar et al. (2022), which relied on a learned objectness score to relabel the background as potential unknowns. Another direction of research focused on clustering classes to better isolate unknowns (Wu et al., 2022b; Yu et al., 2022). Additionally, some works, like Kim et al. (2022); Wu et al. (2022a), introduced a decoupled approach where the classification and localization heads are separated. The objective is to remove label information and instead capture the shared features that makes these labelled items relevant as objects.
@@
However, these works fail to address a crucial problem, which is defining what constitutes an "unknown" object. Currently, there is no clear definition or prior knowledge available to effectively distinguish unknowns from the background. Its interpretation greatly varies depending on the context. For example, in a driving scene, a "debris on the road" could be considered an unknown object  Balasubramaniam and Pasricha (2022), while in a camera surveillance context, it might be perceived as part of the background  Ingle and Kim (2022). Without considering the context, these works can only learn to differentiate knowns and unknowns at low level features such as texture or shape. As a consequence, they miss to model any hierarchical structures and similarities between known and unknown items, whether at the image level or dataset level.
@@
Acknowledging this context information, we argue that a hierarchical structural relationship must exist between the objects to be discovered and the known items  Hosoya et al. (2022). This hierarchy is characterized by classes that share the same semantic context, belonging to the same category such as vehicles, animals, or electronics. Such hierarchical relationships enable the retrieval of common features and facilitate the discovery of unknown objects. For instance, a model trained on objects related to driving scenes can adequately detect stop signs or traffic lights but is not expected to recognize unrelated objects like a couch or any furniture.
@@
In light of this discrepancy, we propose to learn hierarchical relationships between items in order to effectively utilize the representation of known objects for the discovery of unknown items. Ideally, items belonging to the same family (or category) should be closer to each other while being further away from different families (e.g., animals versus vehicles). To capture these structures, Hyperbolic Distance (Nickel and Kiela, 2018; Park et al., 2021), which naturally maps hierarchical latent structures, such as graphs or trees, emerges as a natural distance metric. This mapping in the hyperbolic space exhibits the desirable property of capturing the affinity between unknown items and known items, thereby enhancing the detection of unknown objects.
@@
Figure 1: t-SNE plot of the learned class representations, with colors representing their respective categories. Our SuperClass Regularizer (right) enhances the hierarchical structure by grouping together classes from the same category while pushing apart classes from different categories.
@@
Motivated by the aforementioned literature gap, we propose a Hyperbolic Distance-based Adaptive Relabeling Scheme for Open World Object Detection (dubbed Hyp-OW). Our contribution can be summarized in three parts:
  • Hyp-OW is a simple yet effective method that learns inherent hierarchical structure between objects grouping item from the same category closer while pushing classes from different categories further apart through a SuperClass Regularizer (illustrated in Figure 1, right).
@@
  • Our experiments demonstrate significant improvements in both unknown detection (up to $6\%$) and known accuracy performance (up to $5\%$) with Hyp-OW . These gains are particularly prominent when evaluating on our (designed) Hierarchical dataset, highlighting the advantages of our method in the presence of high inherent hierarchical structures.
2 Related Work
2.1 Open World Object Detection
@@
The OWOD framework, introduced by Joseph et al. (2021), has inspired many recent works due to its realistic and close-to-real-world setting that integrates newly discovered items into the base knowledge progressively. While the first stream of work was originally based on the Faster-RCNN model (Joseph et al., 2021; Yu et al., 2022; Wu et al., 2022b, a), more recent works have utilized Deformable Transformers due to their superior performance (Gupta et al., 2022; Zohar et al., 2022). Joseph et al. (2021) introduced ORE, a Faster-RCNN-based model that learns class prototypes using contrastive learning with Euclidean distance. However, their approach relied on a held-out validation set where unknown items are explicitly labeled to learn an energy-based model to discriminate unknown items. Yu et al. (2022) extended this setting by minimizing the overlap between the distributions of unknown and known classes. OW-DETR (Gupta et al., 2022) designed a novelty-branch head to relabel the top-k highest background scores as unknowns. These pseudo-labels relied on unmatched bounding box proposals with high backbone activation being selected as unknown objects. On the other hand, Wu et al. (2022a) decoupled the localization and classification tasks (introduced by (Kim et al., 2022)) by learning a class-free head to localize objects. Recently, PROB (Zohar et al., 2022) learned a probabilistic objectness score by learning common statistics for all objects using Mahalanobis distance (Lee et al., 2018) and considered all the remaining bounding box proposals as unknown items. During the evaluation phase, they filter out proposal bounding boxes using the latter probabilistic models.
2.2 Class-Agnostic Object Detection
@@
Another stream of work in the field of object detection is dubbed class-agnostic object detection, which focuses on localizing objects  (Kim et al., 2022; Wu et al., 2022a; Jaiswal et al., 2021). The objective is to remove the class label information and learn a shared low-level feature representation that effectively captures the essence of an object.  Kim et al. (2022) designed a pure localization head by introducing a second branch that is decoupled from the classification head.  Jaiswal et al. (2021) introduced an adversarial objective loss function that penalizes label information in the encoded features. Pixel-wise class-free object detection  Gonçalves et al. (2022) used texture gray level quantization to retrieve objects. Saito et al. Saito et al. (2022) designed a new data augmentation method that pastes an annotated object onto an object-free background.  Maaz et al. (2022) leveraged language models to improve unknown detection with their Multi-Modal Vision Transformers.
2.3 Learning hierarchical Representation with Hyperbolic Distance
@@
Pointcare embeddings have been widely used in the literature to learn hierarchical structures from complex symbolic or multi-relational data, which can be represented by graphs or trees, such as social networks or taxonomies  (Nickel and Kiela, 2018; Law et al., 2019). Due to its good performance, it has been applied to image classification as well  (Khrulkov et al., 2020a; Yan et al., 2021; Yue et al., 2023; Ermolov et al., 2022). For example,  Yan et al. (2021) used hierarchical clustering to approximate a multi-layered tree structure representation that guides the hyperbolic distance learning process. Similarly, Liu et al. (2020) used taxonomy embedding from GloVe  (Pennington et al., 2014) to learn a finer-grained representation. Hyperbolic distance has also been used for object detection  (Lang et al., 2022; Ge et al., 2022).  Ge et al. (2022) was interested in learning context-object association rules by reasoning on different image scales. However, none of them leveraged the learned hyperbolic distance to retrieve unknowns items for OWOD.
3 Background
3.1 Problem Formulation
@@
OWOD framework describes the setting where a user receives over the time a stream of $T$ tasks indexed by $t\in[1,T]$. Every task $t$ contains $C_{t}\in\mathbb{N}^{*}$ known classes ( denoted by set $\mathcal{K}^{t}$111whenever there is no ambiguity we will remove the task index $t$ to de-clutter the notation). The goal is to train an object detector module $f$ to accurately recognize the known classes but also discovering unknown classes (denoted by set $\mathcal{U}^{t}$). At the end of task $t$, $\Delta_{t}$ unknown classes are labelled (with an oracle) and included in the next task $t+1$ (containing now $\sum_{j=1}^{t}\Delta_{j}=C_{t}$ known classes). The process repeat until task $T$ that does not contain anymore unknowns.
@@
The dataset of task $t$ is defined as $\mathcal{D}^{t}=\{\mathcal{I}^{t},\mathcal{Y}^{t}\}$ where $\mathcal{I}^{t}=\{\mathbf{I}_{t,1},\mathbf{I}_{t,2},...,\mathbf{I}_{t,N_{t}}\}$, are image inputs and $\mathcal{Y}^{t}=\{\mathbf{y}_{t,1},\mathbf{y}_{t,2},...,\mathbf{y}_{t,N_{t}}\}$ are the corresponding labels (There are $N_{t}$ images for task $t$). Each image $I_{t,i}$ contains a set of annotations $\mathbf{Y}_{t,i}=[\mathbf{l}_{t,i},\mathbf{x}_{t,i},\mathbf{y}_{t,i},\mathbf{w}_{t,i},\mathbf{h}_{t,i}]$ where $\mathbf{l}_{t,i}\in\{0,1\}^{C}$ denotes the object classes and $[\mathbf{x}_{t,i},\mathbf{y}_{t,i},\mathbf{w}_{t,i},\mathbf{h}_{t,i}]$ are the bounding box coordinates. Throughout the training, we will be storing item in a replay buffer $\mathcal{M}$ with a capacity of $m$ exemplar per class. We denote $\mathcal{B}$ the incoming batch.
We follow the setting of OWOD Joseph et al. (2021) where a set of $K$ examplars of each class is stored in a replay buffer at the end of each task $t$ (to mitigate forgetting). Those examplars are then replayed, i.e after task $2$, we replay $(2\cdot K\cdot 20)$ examplars both coming from task $1$ and $2$ ($K$ examplars from each task and $20$ classes per task).
3.2 Deformable Transformers for OWOD
@@
We adopt Deformable Transformers (Zhu et al., 2021) as our base detector, as suggested by Gupta et al. (2022), due to its simplicity and high performance. This architecture processes the image input through a set of encoder-decoder modules to produce $Q$ queries output embeddings $\{\mathbf{q_{i}}\}_{i=1}^{Q}$, where $\mathbf{q_{i}}\in\mathbb{R}^{d}$ ($d$ being the query embedding dimension). These queries are then passed to the bounding box and classification heads, which respectively localize the labeled items and predict their classes. A bipartite matching algorithm (specifically, the Hungarian algorithm (Kuhn, 1955)) is used to match the labeled ground-truth items with each query. The remaining challenge is then to determine whether unmatched queries contain potential unknowns.
3.3 Hyperbolic Embeddings
@@
A Hyperbolic space is a $n$-dimensional Riemann manifold defined as $(\mathbb{B}_{c}^{n},g^{\mathbb{M}})$ with its Poincare ball $\mathbb{B}_{c}^{n}=\{x\in\mathbb{R}^{n}:c\lVert x\rVert^{2}\leq 1,c\geq 0\}$ ($c$ being the constant curvature) and equipped with a Riemannian metric $g^{\mathbb{M}}=\lambda_{x}^{2}g^{E}$ where $\lambda_{x}^{c}=\frac{2}{1-c\lVert x\rVert^{2}}$, $g^{E}=\mathbf{I}_{n}$ is the Euclidian metric tensor. The transformation from the Euclidian to hyperbolic space is done via a bijection termed exponential mapping $\exp_{\mathbf{b}}^{c}:\mathbb{R}^{n}\rightarrow\mathbb{B}_{c}^{n}$.
@@
$\exp_{\mathbf{b}}^{c}(\mathbf{x})=\mathbf{b}\otimes_{c}(\tanh{(\sqrt{c}\frac{\lambda_{\mathbf{b}}^{c}\lVert\mathbf{x}\rVert}{2})}\frac{\lVert\mathbf{x}\rVert}{\sqrt{c}\lVert\mathbf{x}\rVert})$ (1)
@@
where $\mathbf{b}$ represents the base point. The latter is often empirically taken as $\mathbf{b}=\mathbf{0}$ to simplify the formulas without impacting much the results (Ermolov et al., 2022). We will also adopt this value in our study.
@@
$d_{hyp}(\mathbf{x},\mathbf{y})=\frac{2}{\sqrt{c}}\arctan{(\sqrt{c}\lVert-\mathbf{x}\otimes_{c}\mathbf{y}\rVert)}$ (2)
where the addition operation $\otimes_{c}$ is defined as : $\mathbf{x}\otimes_{c}\mathbf{y}=\frac{(1+2c\left<\mathbf{x},\mathbf{y}\right>+c\lVert\mathbf{y}\rVert^{2})\mathbf{x}+(1-c\lVert\mathbf{x}\rVert^{2})\mathbf{y}}{1+2c\left<\mathbf{x},\mathbf{y}\right>+c^{2}\lVert\mathbf{x}\rVert^{2}\lVert\mathbf{y}\rVert^{2}}$.
@@
From now on, we will denote $\mathbf{z_{i}}$ the projection of the queries $\mathbf{q_{i}}$ into the hyperbolic embedding space, i.e, $\mathbf{z_{i}}=\exp_{\mathbf{b}}^{c}(\mathbf{q_{i}})$. Note that when $c=0$, the distance boils down to the cosine similarity defined as:
$D_{cos}(\mathbf{x},\mathbf{y})=2-2\frac{<\mathbf{x},\mathbf{y}>}{\lVert\mathbf{x}\rVert_{2}\cdot\lVert\mathbf{y}\rVert_{2}}$ (in this case no exponential mapping is needed, i.e $\mathbf{z_{i}}=\mathbf{q_{i}}$222from now on, we will refer to $\mathbf{z_{i}}$ interchangeably for both the query and its corresponding embedding $\mathbf{q_{i}}$)
4 Hyp-OW
@@
In this section, we provide a detailed explanation of each module of our proposed method. Hyp-OW can be summarized by three main components (Figure 2): a Hyperbolic Metric Distance learning (Section 4.1), a SuperClass Regularizer (Section 4.2) and an Adaptive Relabeling Scheme to detect unknowns (Section 4.3).
@@
Figure 2: Overview of each component of Hyp-OW.There are three main components: the Hyperbolic Contrastive Loss, which learns a hierarchical structural representation of each class; the SuperClass Regularizer, which models the semantic relationship among classes to ensure proximity within the same category and distance from different categories; and the Adaptive Relabeling module, which utilizes the hierarchical structure to compute the hyperbolic distance between candidate proposals and known items. If this distance $d$ is lower than a certain threshold ($\delta$), the proposal is relabelled as unknown.
4.1 Metric Learning with Hyperbolic Distance
@@
We learn feature representation in the hyperbolic embedding space using a contrastive loss. The idea is to move closer features belonging to the same class $c$ while repetlling then away from any features from class $j\neq c$. Let’s denote $\mathbf{z_{i}^{c}}$ any query $i$ matched with class $c\in\mathcal{K}$333by abuse of notation, we will also use $c$ for the class label since the meaning can be infered from the context. During the training, we maintain a replay buffer $\mathcal{M}$ where we store $m$ embedding features per class.
@@
For every query element $\mathbf{z_{i}^{c}}$ of the incoming batch $\mathcal{B}$, we sample $k=1$ elements of the same class $c\in\mathcal{K}$ from the replay buffer $\mathcal{M}$ (which will be considered as the positive examples, the $22|\mathcal{B}|-$ 2 remaining samples be considered as the negative examples).
For simplicity we denote all the elements from the buffer and the batch $\mathcal{A}=\mathcal{B}\cup\mathcal{M}$. Every element $\mathbf{z_{i}^{c}}\in\mathcal{B}$ (respectively $i\in\mathcal{M}$) has its positive counterpart examples $\mathbf{z_{i^{+}}}$ of same class $c$ from $\mathcal{M}$ (respectively from $\mathcal{B}$) and $|\mathcal{A}|-2$ negative examples (both from $\mathcal{B}$ and $\mathcal{M}$) denoted as $\mathbf{z_{i^{-}}}$.
@@
Defining a temperature $\tau_{1}$, the contrastive loss is then expressed as:
$\mathscr{L}_{hyp}=-\displaystyle{\sum_{i\in\mathcal{A},c\in\mathcal{K}}}\log\frac{\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{z_{i^{+}}})}{\tau_{1}})}{\displaystyle{\sum_{i^{-}\in\mathcal{A}\setminus\{i,i^{+}\}}}\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{z_{i^{-}}})}{\tau_{1}})}$ (3)
@@
This loss aims at attracting representation of $\mathbf{z_{i}^{c}}$ closer to its positive counterpart $\mathbf{z_{i^{+}}}$ while repelling from other classes representation $\mathbf{z_{i^{-}}}$,$i\in\mathcal{A},c\in\mathcal{K}$.
4.2 SuperClass Regularization
Many real-world datasets exhibit a natural hierarchical structure, where classes can be organized into categories. For instance, dogs and cats belong to the animal category, while cars and trucks belong to vehicles category. To leverage this hierarchical information, we propose a SuperClass Regularizer (we will use SuperClass and category interchangeably throughout this study), which encourages classes within the same category to be closer in the embedding space while pushing them away from classes in different categories.

### v2

@@ Abstract
Open World Object Detection (OWOD) is a challenging and realistic task that extends beyond the scope of standard Object Detection task. It involves detecting both known and unknown objects while integrating learned knowledge for future tasks. However, the level of ”unknownness” varies significantly depending on the context. For example, a tree is typically considered part of the background in a self-driving scene, but it may be significant in a household context. We argue that this contextual information should already be embedded within the known classes. In other words, there should be a semantic or latent structure relationship between the known and unknown items to be discovered. Motivated by this observation, we propose Hyp-OW, a method that learns and models hierarchical representation of known items through a SuperClass Regularizer. Leveraging this representation allows us to effectively detect unknown objects using a similarity distance-based relabeling module. Extensive experiments on benchmark datasets demonstrate the effectiveness of Hyp-OW, achieving improvement in both known and unknown detection (up to 6 percent). These findings are particularly pronounced in our newly designed benchmark, where a strong hierarchical structure exists between known and unknown objects. Our code can be found at https://github.com/tldoan/-HYP-OW-AAAI-2024-.
Introduction
@@
Advances in Object Detection (OD) have unlocked a plethora of practical applications such as robotics  (Zhou et al. 2022), self-driving cars  (Balasubramaniam and Pasricha 2022), manufacturing  (Malburg et al. 2021), and medical analysis  (Yang and Yu 2021). Recent breakthroughs in attention-based neural network architecture, such as Deformable Transformers  (Zhu et al. 2021), have yielded impressive performance in these settings. However, most of these approaches assume a fixed number of classes (closed-world assumption), which is rare in reality. Continual Object Detection  (Menezes et al. 2023) takes a step further by incrementally adding new classes, resulting in a distribution shift in the input and the well-known phenomenon of catastrophic forgetting (Kirkpatrick et al. 2017; Doan et al. 2021) where the network forgets previously learned knowledge. Open World (OW) (Bendale and Boult 2015) takes these assumptions even further, introducing the detection and integration of newly discovered classes.
@@
While the seminal work by  (Bendale and Boult 2015) introduced OW framework, further advancements by  (Joseph et al. 2021) extended it in two key aspects: the detection task and continual learning. However, a significant challenge within this framework lies in the absence of annotations for unknown objects, leading to biases toward known labels and potential confusion between unknown items and the background. This bias significantly impedes the accurate identification of unknown objects and presents a major hurdle in the detection process.
@@
Previous approaches, often relying on shared features or objectness scores (Joseph et al. 2021; Gupta et al. 2022; Zohar, Wang, and Yeung 2023), as well as clustering methods (Wu et al. 2022b; Yu et al. 2022), have failed to address a critical challenge: defining what constitutes an ”unknown” object. Currently, there is no clear definition or prior knowledge available to effectively distinguish unknowns from the background. Its interpretation greatly varies depending on the context. For example, in a driving scene, a ”debris on the road” could be considered an unknown object  (Balasubramaniam and Pasricha 2022), while in a camera surveillance context, it might be perceived as part of the background  (Ingle and Kim 2022). Without considering the context, these works can only learn to differentiate knowns and unknowns at low level features such as texture or shape. As a consequence, they fail to model any hierarchical structures and similarities between known and unknown items, whether at the image level or dataset level.
@@
Acknowledging this context information, we argue that a hierarchical structure must exist between the objects to be discovered and the known items  (Hosoya, Suganuma, and Okatani 2022). This hierarchy is characterized by classes that share the same semantic context, belonging to the same category such as vehicles, animals, or electronics. Such hierarchical relationships enable the retrieval of common features and facilitate the discovery of unknown objects. For instance, a model trained on objects related to driving scenes can adequately detect stop signs or traffic lights but is not expected to recognize unrelated objects like a couch or any furniture.
@@
Given this discrepancy, we propose modeling hierarchical relationships among items to enhance the discovery of unknowns. Ideally, items belonging to the same family (or category) should be closer to each other while being further away from different families (e.g., animals versus vehicles). To capture these structures, Hyperbolic Distance (Nickel and Kiela 2018; Park et al. 2021), which naturally maps hierarchical latent structures, such as graphs or trees, emerges as an ideal distance metric. This has the desirable property of capturing the affinity between unknown items and known items, thereby enhancing the detection of unknown objects.
@@
Figure 1: t-SNE plot of the learned class representations, with colors representing their respective categories. Our SuperClass Regularizer (right) learns the hierarchical structure by grouping together classes from the same category while pushing apart those from different categories.
@@
Motivated by the aforementioned literature gap, we propose a Hyperbolic Distance-based Adaptive Relabeling Scheme for Open World Object Detection (Hyp-OW). Our contribution can be summarized in three parts:
  • Hyp-OW is a simple yet effective method that learns innate hierarchical structure between objects grouping item from the same category closer while pushing classes from different categories further apart through a SuperClass Regularizer (illustrated in Figure 1, right).
@@
  • Our experiments demonstrate significant improvements in both unknown recall (up to $6\%$) and known object detection performance (up to $5\%$) with Hyp-OW . These gains are particularly prominent when evaluating on our (designed) Hierarchical dataset that exhibits high inherent hierarchical structures.
Related Work
Open World Object Detection
@@
The OWOD framework, introduced by (Joseph et al. 2021), has inspired many recent works due to its realistic and close-to-real-world setting that integrates newly discovered items into the base knowledge progressively. While the first stream of work was originally based on the Faster-RCNN model (Joseph et al. 2021; Yu et al. 2022; Wu et al. 2022b, a), more recent works have utilized Deformable Transformers due to their superior performance (Gupta et al. 2022; Zohar, Wang, and Yeung 2023). Joseph et al. (2021) introduced ORE, a Faster-RCNN-based model that learns class prototypes using contrastive learning with Euclidean distance. However, their approach relied on a held-out validation set where unknown items are explicitly labeled to learn an energy-based model to discriminate unknown items. (Yu et al. 2022) extended this setting by minimizing the overlap between the distributions of unknown and known classes. OW-DETR (Gupta et al. 2022) designed a novelty-branch head to relabel the top-k highest background scores as unknowns. These pseudo-labels relied on unmatched bounding box proposals with high backbone activation being selected as unknown objects. On the other hand, Wu et al. (2022a) decoupled the localization and classification tasks (introduced by Kim et al. (2022)) by learning a class-free head to localize objects. Recently, PROB (Zohar, Wang, and Yeung 2023) learned a probabilistic objectness score by learning common statistics for all objects using Mahalanobis distance (Lee et al. 2018) and considered all the remaining bounding box proposals as unknown items. During the evaluation phase, they filter out proposal bounding boxes using the latter probabilistic models.
Class-Agnostic Object Detection
@@
Another stream of work in the field of object detection is dubbed class-agnostic object detection, which focuses on localizing objects  (Kim et al. 2022; Wu et al. 2022a; Jaiswal et al. 2021). The objective is to remove the class label information and learn a shared low-level feature representation that effectively captures the essence of an object.  Kim et al. (2022) designed a pure localization head by introducing a second branch that is decoupled from the classification head.  Jaiswal et al. (2021) introduced an adversarial objective loss function that penalizes label information in the encoded features. Pixel-wise class-free object detection  (Gonçalves et al. 2022) used texture gray level quantization to retrieve objects. Saito et al. Saito et al. (2022) designed a new data augmentation method that pastes an annotated object onto an object-free background.  Maaz et al. (2022) leveraged language models to improve unknown detection with their Multi-Modal Vision Transformers.
Learning Hierarchical Representation with Hyperbolic Distance
@@
Poincare embeddings have been widely used in the literature to learn hierarchical structures from complex symbolic or multi-relational data, which can be represented by graphs or trees, such as social networks or taxonomies  (Nickel and Kiela 2018; Law et al. 2019). Due to its good performance, it has been applied to image classification as well  (Khrulkov et al. 2020a; Yan et al. 2021; Yue et al. 2023; Ermolov et al. 2022). For example,  Yan et al. (2021) used hierarchical clustering to approximate a multi-layered tree structure representation that guides the hyperbolic distance learning process. Similarly, Liu et al. (2020) used taxonomy embedding from GloVe  (Pennington, Socher, and Manning 2014) to learn a finer-grained representation. Hyperbolic distance has also been used for object detection  (Lang et al. 2022; Ge et al. 2022).  Ge et al. (2022) was interested in learning context-object association rules by reasoning on different image scales. However, none of them leveraged the learned hyperbolic distance to retrieve unknowns items for OWOD.
Background
Problem Formulation
@@
OWOD framework describes the setting where a user receives over the time a stream of $T$ tasks indexed by $t\in[1,T]$. Every task $t$ contains $C_{t}\in\mathbb{N}^{*}$ known classes (denoted by set $\mathcal{K}^{t}$ [^1] Whenever there is no ambiguity we will remove the task index $t$ to de-clutter the notation.). The goal is to train an object detector module $f$ to accurately recognize the known classes but also discovering unknown classes (denoted by set $\mathcal{U}^{t}$). At the end of task $t$, $C_{t+1}$ unknown classes are labelled (with an oracle) and included in the next task $t+1$. The process repeat until task $T$ that does not contain anymore unknowns.
@@
The dataset of task $t$ is defined as $\mathcal{D}^{t}=\{\mathcal{I}^{t},\mathcal{Y}^{t}\}$ where $\mathcal{I}^{t}$ are image inputs and $\mathcal{Y}^{t}$ the corresponding labels. Each label consists of a list of bounding box locations along with their corresponding labels. We follow the setting of OWOD (Joseph et al. 2021) where a set of $K$ examplars of each class is stored in a replay buffer at the end of each task $t$ (to mitigate forgetting) to be replayed. Additionally, throughout the training, we will be storing item in a replay buffer $\mathcal{M}$ with a capacity of $m$ exemplar per class. We denote $\mathcal{B}$ the incoming batch.
Deformable Transformers for OWOD
@@
We adopt Deformable Transformers (Zhu et al. 2021) as our base detector, as it showed simplicity and high performance (Gupta et al. 2022). The image input is processed through a set of encoder-decoder modules to output $Q$ queries $\{\mathbf{q_{i}}\}_{i=1}^{Q}$, where $\mathbf{q_{i}}\in\mathbb{R}^{d}$ are the output embeddings. These queries served as input to different heads such as classification and localization heads. since $Q$ is higher than the number of ground-truth labels, the Hungarian algorithm (Kuhn 1955) is used to match the labeled ground-truth items with each query. We refer the reader to (Zhu et al. 2021) for more details.
Hyperbolic Embeddings
@@
A Hyperbolic space is a $n$-dimensional Riemann manifold defined as $(\mathbb{B}_{c}^{n},g^{\mathbb{M}})$ with its Poincare ball $\mathbb{B}_{c}^{n}=\{x\in\mathbb{R}^{n}:c\lVert x\rVert^{2}\leq 1,c\geq 0\}$ ($c$ being the constant curvature) and equipped with a Riemannian metric $g^{\mathbb{M}}=(\frac{2}{1-\lVert\mathbf{x}\rVert^{2}})^{2}g^{E}$ where $g^{E}=\mathbf{I}_{n}$ is the Euclidian metric tensor. The transformation from the Euclidian to hyperbolic space is done via a bijection termed exponential mapping $\exp_{\mathbf{b}}^{c}:\mathbb{R}^{n}\rightarrow\mathbb{B}_{c}^{n}$.
@@
$\exp_{\mathbf{b}}^{c}(\mathbf{x})=\mathbf{b}\oplus_{c}(\tanh{(\sqrt{c}\frac{\lambda_{\mathbf{b}}^{c}\lVert\mathbf{x}\rVert}{2})}\frac{\lVert\mathbf{x}\rVert}{\sqrt{c}\lVert\mathbf{x}\rVert})$ (1)
@@
with $\lambda_{\mathbf{b}}^{c}=\frac{2}{1-c\lVert\mathbf{b}\rVert^{2}}$ and the base point $\mathbf{b}$. The latter is often empirically taken as $\mathbf{b}=\mathbf{0}$ to simplify the formulas without impacting much the results (Ermolov et al. 2022). We will also adopt this value in our study.
@@
$d_{hyp}(\mathbf{x},\mathbf{y})=\frac{2}{\sqrt{c}}\arctan{(\sqrt{c}\lVert-\mathbf{x}\oplus_{c}\mathbf{y}\rVert)}$ (2)
where the addition operation $\oplus_{c}$ is defined as :
$\mathbf{x}\oplus_{c}\mathbf{y}=\frac{(1+2c\left<\mathbf{x},\mathbf{y}\right>+c\lVert\mathbf{y}\rVert^{2})\mathbf{x}+(1-c\lVert\mathbf{x}\rVert^{2})\mathbf{y}}{1+2c\left<\mathbf{x},\mathbf{y}\right>+c^{2}\lVert\mathbf{x}\rVert^{2}\lVert\mathbf{y}\rVert^{2}}$.
@@
From now on, we will denote $\mathbf{z_{i}}$ the projection of the queries $\mathbf{q_{i}}$ into the hyperbolic embedding space, i.e, $\mathbf{z_{i}}=\exp_{\mathbf{b}}^{c}(\mathbf{q_{i}})$. When $c\to 0$, we recover the Euclidian distance: $\lim_{c\to 0}d_{hyp}(\mathbf{x},\mathbf{y})=2\rVert\mathbf{x}-\mathbf{y}\rVert$. This quantity is also related to the cosine similarity $d_{cos}(\mathbf{x},\mathbf{y})=2-2\frac{<\mathbf{x},\mathbf{y}>}{\lVert\mathbf{x}\rVert\cdot\lVert\mathbf{y}\rVert}$ in the case of normalized vectors (See supplementary).
Hyp-OW
@@
In this section, we provide a detailed explanation of each module of our proposed method. Hyp-OW can be summarized by three main components (Figure 2): a Hyperbolic Metric Distance learning, a SuperClass Regularizer, and an Adaptive Relabeling Scheme to detect unknowns.
@@
Figure 2: Overview of Hyp-OW. Comprising three core components: the Hyperbolic Contrastive Loss for representation learning at the class level; the SuperClass Regularizer, for semantic relationships at the category level; and the Adaptive Relabeling module, for unknown retrieval with the previously learned representation. If a distance $d$ between a candidate proposal and known items is lower than a certain threshold ($\delta$), the proposal is relabelled as unknown.
Metric Learning with Hyperbolic Distance
@@
We learn feature representation in the hyperbolic embedding space using a contrastive loss. The idea is to move closer features belonging to the same class $c$ [^2] By abuse of notation, we will also use $c$ for the class label since the meaning can be inferred from the context while repelling them from features of different classes. Let’s denote $\mathbf{z_{i}^{c}}$ any query $i$ matched with class $c\in\mathcal{K}$. To facilitate readability, we will omit the index $c$ whenever there is no confusion about the class context.
@@
Throughout training, we maintain a replay buffer $\mathcal{M}$ where we store $m$ embedding features per class. For every query element $\mathbf{z_{i}}$ of the incoming batch $\mathcal{B}$, we sample $k=1$ element of the same class from the replay buffer $\mathcal{M}$ denoted $\mathbf{z_{i^{+}}}$ (this element serves as the positive comparison) and consider the $22|\mathcal{B}|-$ 2 remaining samples as the negative examples $\mathbf{z_{i^{-}}}$.
@@
If we denote $\mathcal{A}=\mathcal{B}\cup\mathcal{M}$ and define a temperature $\tau_{1}$, the contrastive loss is then expressed as:
$\mathscr{L}_{hyp}=-\displaystyle{\sum_{i\in\mathcal{A}}}\log\frac{\exp(\frac{-d_{hyp}(\mathbf{z_{i}},\mathbf{z_{i^{+}}})}{\tau_{1}})}{\displaystyle{\sum_{i^{-}\in\mathcal{A}\setminus\{i,i^{+}\}}}\exp(\frac{-d_{hyp}(\mathbf{z_{i}},\mathbf{z_{i^{-}}})}{\tau_{1}})}$ (3)
@@
This loss aims at attracting representation of $\mathbf{z_{i}}$ closer to its positive counterpart $\mathbf{z_{i^{+}}}$ while repelling from the negative examples $\mathbf{z_{i^{-}}}$, $i\in\mathcal{A}$.
SuperClass Regularization
Numerous real-world datasets inherently possess hierarchical structures, allowing classes to be categorized. For instance, dogs and cats fall under the broader category of ”animals,” while cars and trucks belong to the category of ”vehicles.” To harness this inherent hierarchy, we introduce a SuperClass Regularizer (we will use ”SuperClass” and ”category” interchangeably in this context). In contrast to Eq 3, our proposed regularization encourages grouping at the SuperClass level rather than the class level.